# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v1

@@ 6 DISCUSSION
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0035 0.0022 0.0050 0.0102
1.2 0.0072 0.0063 0.0056 0.0171
1.3 0.0081 0.0088 0.0053 0.0172
1.4 0.0074 0.0077 0.0061 0.0168
1.5 0.0081 0.0079 0.0059 0.0149
1.7 0.0057 0.0057 0.0050 0.0133
1.9 0.0049 0.0047 0.0039 0.0108
2.0 0.0053 0.0056 0.0046 0.0103
2.5 0.0045 0.0046 0.0043 0.0106
3.0 0.0031 0.0029 0.0034 0.0086
Case 1b Weak correlation ZI 1.1 0.0010 0.0017 0.0060 0.0197
1.2 0.0044 0.0034 0.0096 0.0214
1.3 0.0067 0.0070 0.0070 0.0156
1.4 0.0086 0.0085 0.0051 0.0157
1.5 0.0089 0.0092 0.0059 0.0138
1.7 0.0067 0.0065 0.0056 0.0136
1.9 0.0058 0.0057 0.0048 0.0117
2.0 0.0049 0.0050 0.0045 0.0118
2.5 0.0032 0.0034 0.0032 0.0089
3.0 0.0036 0.0037 0.0042 0.0095
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0010 0.0065 0.0193
1.2 0.0107 0.0128 0.0063 0.0150
1.3 0.0114 0.0132 0.0069 0.0160
1.4 0.0120 0.0134 0.0060 0.0146
1.5 0.0095 0.0098 0.0069 0.0157
1.7 0.0072 0.0079 0.0059 0.0145
1.9 0.0052 0.0056 0.0058 0.0128
2.0 0.0042 0.0045 0.0048 0.0125
2.5 0.0035 0.0035 0.0041 0.0102
3.0 0.0033 0.0032 0.0032 0.0093
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0055 0.0122
1.2 0.0093 0.0106 0.0065 0.0186
1.3 0.0117 0.0130 0.0069 0.0168
1.4 0.0102 0.0114 0.0061 0.0153
1.5 0.0066 0.0076 0.0054 0.0149
1.7 0.0053 0.0057 0.0058 0.0134
1.9 0.0039 0.0040 0.0059 0.0126
2.0 0.0047 0.0050 0.0045 0.0110
2.5 0.0039 0.0037 0.0042 0.0103
3.0 0.0030 0.0032 0.0041 0.0094
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0073 0.0133
1.2 0.0084 0.0114 0.0074 0.0183
1.3 0.0073 0.0084 0.0080 0.0184
1.4 0.0086 0.0096 0.0069 0.0145
1.5 0.0061 0.0069 0.0067 0.0150
1.7 0.0045 0.0050 0.0066 0.0136
1.9 0.0045 0.0050 0.0051 0.0129
2.0 0.0039 0.0042 0.0049 0.0118
2.5 0.0033 0.0035 0.0037 0.0096
3.0 0.0030 0.0029 0.0031 0.0085
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0040 0.0171
1.2 0.0106 0.0120 0.0084 0.0162
1.3 0.0077 0.0090 0.0063 0.0161
1.4 0.0078 0.0083 0.0061 0.0153
1.5 0.0068 0.0073 0.0051 0.0125
1.7 0.0044 0.0049 0.0049 0.0119
1.9 0.0030 0.0032 0.0050 0.0112
2.0 0.0035 0.0038 0.0052 0.0113
2.5 0.0033 0.0034 0.0040 0.0092
3.0 0.0036 0.0037 0.0039 0.0094
@@
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0016 0.0014 0.0009 0.0021
1.2 0.0427 0.0422 0.0346 0.0435
1.3 0.1227 0.1243 0.0899 0.1041
1.4 0.1923 0.1942 0.1432 0.1606
1.5 0.2444 0.2472 0.1915 0.2115
1.7 0.3153 0.3180 0.2685 0.2897
1.9 0.3692 0.3707 0.3258 0.3475
2.0 0.3926 0.3957 0.3495 0.3720
2.5 0.4706 0.4729 0.4346 0.4554
3.0 0.5251 0.5273 0.4920 0.5128
Case 1b Weak correlation ZI 1.1 0.0012 0.0012 0.0011 0.0022
1.2 0.0406 0.0399 0.0331 0.0416
1.3 0.1216 0.1232 0.0848 0.1006
1.4 0.1948 0.1968 0.1401 0.1581
1.5 0.2426 0.2451 0.1888 0.2102
1.7 0.3182 0.3196 0.2675 0.2898
1.9 0.3709 0.3733 0.3261 0.3457
2.0 0.3921 0.3939 0.3467 0.3678
2.5 0.4717 0.4742 0.4327 0.4564
3.0 0.5227 0.5253 0.4910 0.5124
Case 2a Moderate correlation No ZI 1.1 0.0061 0.0060 0.0012 0.0026
1.2 0.2096 0.2170 0.0526 0.0674
1.3 0.3304 0.3377 0.1334 0.1530
1.4 0.3991 0.4036 0.1998 0.2210
1.5 0.4489 0.4537 0.2499 0.2751
1.7 0.5172 0.5197 0.3359 0.3637
1.9 0.5663 0.5681 0.4043 0.4323
2.0 0.5837 0.5847 0.4310 0.4589
2.5 0.6396 0.6411 0.5363 0.5608
3.0 0.6734 0.6746 0.5984 0.6214
Case 2b Moderate correlation ZI 1.1 0.0067 0.0072 0.0011 0.0025
1.2 0.2067 0.2133 0.0523 0.0659
1.3 0.3309 0.3367 0.1348 0.1535
1.4 0.3968 0.4021 0.1973 0.2186
1.5 0.4438 0.4477 0.2475 0.2723
1.7 0.5138 0.5157 0.3340 0.3599
1.9 0.5601 0.5620 0.4000 0.4274
2.0 0.5772 0.5790 0.4297 0.4570
2.5 0.6329 0.6339 0.5313 0.5561
3.0 0.6696 0.6708 0.5967 0.6205
Case 3a Strong correlation No ZI 1.1 0.0065 0.0076 0.0009 0.0020
1.2 0.2680 0.2784 0.0329 0.0455
1.3 0.4595 0.4682 0.1088 0.1313
1.4 0.5675 0.5746 0.1807 0.2100
1.5 0.6272 0.6330 0.2480 0.2791
1.7 0.7015 0.7049 0.3546 0.3868
1.9 0.7460 0.7482 0.4300 0.4612
2.0 0.7629 0.7644 0.4608 0.4940
2.5 0.8096 0.8106 0.5793 0.6094
3.0 0.8330 0.8336 0.6530 0.6799
Case 3b Strong correlation ZI 1.1 0.0083 0.0094 0.0010 0.0021
1.2 0.2627 0.2769 0.0338 0.0471
1.3 0.4588 0.4688 0.1098 0.1326
1.4 0.5638 0.5717 0.1826 0.2115
1.5 0.6275 0.6333 0.2492 0.2801
1.7 0.7044 0.7080 0.3534 0.3848
1.9 0.7485 0.7507 0.4276 0.4599
2.0 0.7630 0.7651 0.4574 0.4893
2.5 0.8100 0.8109 0.5787 0.6097
3.0 0.8341 0.8348 0.6556 0.6825
@@
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0003 0.0001 0.0003
1.4 0.0002 0.0002 0.0001 0.0004
1.5 0.0003 0.0003 0.0002 0.0005
1.7 0.0002 0.0003 0.0002 0.0005
1.9 0.0002 0.0002 0.0002 0.0005
2.0 0.0003 0.0003 0.0002 0.0005
2.5 0.0003 0.0003 0.0002 0.0006
3.0 0.0002 0.0002 0.0002 0.0006
Case 1b Weak correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0001 0.0001 0.0001 0.0002
1.3 0.0002 0.0002 0.0001 0.0002
1.4 0.0003 0.0003 0.0001 0.0004
1.5 0.0003 0.0004 0.0002 0.0004
1.7 0.0003 0.0003 0.0002 0.0005
1.9 0.0003 0.0003 0.0002 0.0005
2.0 0.0002 0.0002 0.0002 0.0006
2.5 0.0002 0.0002 0.0002 0.0005
3.0 0.0003 0.0003 0.0003 0.0006
Case 2a Moderate correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0008 0.0002 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0006 0.0007 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0007
2.0 0.0003 0.0004 0.0003 0.0008
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0002 0.0007
Case 2b Moderate correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0006 0.0001 0.0002
1.3 0.0007 0.0007 0.0001 0.0004
1.4 0.0006 0.0007 0.0002 0.0005
1.5 0.0004 0.0005 0.0002 0.0006
1.7 0.0004 0.0004 0.0003 0.0006
1.9 0.0003 0.0003 0.0003 0.0007
2.0 0.0004 0.0004 0.0003 0.0007
2.5 0.0003 0.0003 0.0003 0.0007
3.0 0.0003 0.0003 0.0003 0.0008
Case 3a Strong correlation No ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0005 0.0007 0.0001 0.0002
1.3 0.0006 0.0007 0.0001 0.0004
1.4 0.0009 0.0010 0.0002 0.0005
1.5 0.0007 0.0007 0.0003 0.0007
1.7 0.0005 0.0006 0.0004 0.0008
1.9 0.0005 0.0006 0.0003 0.0009
2.0 0.0005 0.0005 0.0003 0.0009
2.5 0.0004 0.0004 0.0003 0.0009
3.0 0.0004 0.0004 0.0003 0.0009
Case 3b Strong correlation ZI 1.1 0.0000 0.0000 0.0000 0.0001
1.2 0.0006 0.0007 0.0001 0.0002
1.3 0.0007 0.0008 0.0001 0.0004
1.4 0.0008 0.0009 0.0002 0.0005
1.5 0.0007 0.0008 0.0002 0.0006
1.7 0.0005 0.0006 0.0003 0.0007
1.9 0.0004 0.0004 0.0003 0.0008
2.0 0.0004 0.0005 0.0004 0.0009
2.5 0.0004 0.0004 0.0004 0.0009
3.0 0.0005 0.0005 0.0004 0.0010
@@
Case $\lambda$ DP Poisson local-global DP ZIP local-global DP Hu et al. DP Hu et al. posterior quantile optim
Case 1a Weak correlation No ZI 1.1 0.0030 0.0025 0.0018 0.0040
1.2 0.0783 0.0772 0.0650 0.0810
1.3 0.2129 0.2152 0.1615 0.1844
1.4 0.3180 0.3204 0.2466 0.2722
1.5 0.3887 0.3924 0.3175 0.3444
1.7 0.4761 0.4791 0.4200 0.4446
1.9 0.5358 0.5375 0.4883 0.5115
2.0 0.5602 0.5634 0.5148 0.5380
2.5 0.6369 0.6390 0.6027 0.6213
3.0 0.6859 0.6879 0.6566 0.6738
Case 1b Weak correlation ZI 1.1 0.0023 0.0022 0.0022 0.0042
1.2 0.0748 0.0733 0.0622 0.0774
1.3 0.2109 0.2130 0.1531 0.1787
1.4 0.3212 0.3240 0.2415 0.2682
1.5 0.3863 0.3894 0.3138 0.3429
1.7 0.4793 0.4808 0.4187 0.4449
1.9 0.5378 0.5403 0.4886 0.5094
2.0 0.5602 0.5621 0.5118 0.5333
2.5 0.6382 0.6406 0.6011 0.6226
3.0 0.6837 0.6859 0.6556 0.6731
Case 2a Moderate correlation No ZI 1.1 0.0104 0.0102 0.0024 0.0050
1.2 0.3333 0.3429 0.0973 0.1231
1.3 0.4894 0.4973 0.2322 0.2616
1.4 0.5642 0.5684 0.3299 0.3578
1.5 0.6141 0.6186 0.3966 0.4269
1.7 0.6775 0.6796 0.4997 0.5287
1.9 0.7198 0.7212 0.5724 0.5988
2.0 0.7344 0.7351 0.5992 0.6242
2.5 0.7777 0.7789 0.6954 0.7141
3.0 0.8025 0.8034 0.7463 0.7622
Case 2b Moderate correlation ZI 1.1 0.0114 0.0122 0.0021 0.0049
1.2 0.3292 0.3375 0.0969 0.1204
1.3 0.4900 0.4964 0.2345 0.2626
1.4 0.5624 0.5677 0.3265 0.3548
1.5 0.6101 0.6136 0.3938 0.4237
1.7 0.6750 0.6766 0.4975 0.5248
1.9 0.7150 0.7166 0.5681 0.5941
2.0 0.7289 0.7303 0.5980 0.6226
2.5 0.7727 0.7735 0.6911 0.7102
3.0 0.7999 0.8007 0.7447 0.7614
Case 3a Strong correlation No ZI 1.1 0.0114 0.0131 0.0018 0.0039
1.2 0.4044 0.4167 0.0623 0.0852
1.3 0.6199 0.6280 0.1940 0.2290
1.4 0.7176 0.7231 0.3035 0.3437
1.5 0.7660 0.7701 0.3947 0.4325
1.7 0.8212 0.8234 0.5209 0.5537
1.9 0.8515 0.8527 0.5988 0.6270
2.0 0.8629 0.8638 0.6284 0.6571
2.5 0.8929 0.8934 0.7314 0.7531
3.0 0.9072 0.9076 0.7881 0.8055
Case 3b Strong correlation ZI 1.1 0.0145 0.0162 0.0020 0.0040
1.2 0.3987 0.4159 0.0639 0.0881
1.3 0.6192 0.6287 0.1955 0.2312
1.4 0.7147 0.7211 0.3062 0.3455
1.5 0.7658 0.7702 0.3964 0.4342
1.7 0.8231 0.8255 0.5197 0.5520
1.9 0.8537 0.8551 0.5966 0.6261
2.0 0.8633 0.8645 0.6251 0.6529
2.5 0.8931 0.8936 0.7308 0.7536
3.0 0.9076 0.9080 0.7897 0.8071

### v2

@@ 7 DISCUSSION
Case $\lambda$
DP Poisson
local-global
DP ZIP
local-global
DP Hu et al.
DP Hu et al.
posterior quantile optim
Case 1a
Weak correlation
No ZI
1.1 0.0030 0.0030 0.0040 0.0080
1.2 0.0084 0.0111 0.0075 0.0144
1.3 0.0119 0.0117 0.0074 0.0152
1.4 0.0084 0.0088 0.0040 0.0103
1.5 0.0064 0.0063 0.0037 0.0101
1.7 0.0068 0.0067 0.0040 0.0082
1.9 0.0056 0.0057 0.0038 0.0077
2.0 0.0055 0.0057 0.0030 0.0071
2.5 0.0047 0.0048 0.0029 0.0067
3.0 0.0042 0.0042 0.0026 0.0056
Case 1b
Weak correlation
30% ZI
1.1 0.0010 0.0000 0.0010 0.0053
1.2 0.0079 0.0079 0.0060 0.0114
1.3 0.0084 0.0088 0.0048 0.0102
1.4 0.0074 0.0076 0.0044 0.0103
1.5 0.0063 0.0065 0.0044 0.0098
1.7 0.0058 0.0058 0.0031 0.0089
1.9 0.0056 0.0057 0.0037 0.0080
2.0 0.0055 0.0055 0.0037 0.0078
2.5 0.0046 0.0045 0.0028 0.0065
3.0 0.0037 0.0037 0.0022 0.0049
Case 2a
Moderate correlation
No ZI
1.1 0.0010 0.0000 0.0000 0.0026
1.2 0.0111 0.0128 0.0061 0.0153
1.3 0.0083 0.0082 0.0070 0.0148
1.4 0.0064 0.0067 0.0052 0.0108
1.5 0.0074 0.0072 0.0052 0.0101
1.7 0.0046 0.0047 0.0035 0.0078
1.9 0.0041 0.0041 0.0036 0.0079
2.0 0.0032 0.0032 0.0031 0.0075
2.5 0.0030 0.0030 0.0028 0.0063
3.0 0.0024 0.0023 0.0026 0.0056
Case 2b
Moderate correlation
30% ZI
1.1 0.0033 0.0005 0.0005 0.0082
1.2 0.0124 0.0142 0.0078 0.0159
1.3 0.0105 0.0110 0.0059 0.0133
1.4 0.0056 0.0057 0.0047 0.0121
1.5 0.0056 0.0054 0.0039 0.0098
1.7 0.0045 0.0045 0.0034 0.0088
1.9 0.0033 0.0032 0.0031 0.0073
2.0 0.0027 0.0029 0.0032 0.0073
2.5 0.0025 0.0026 0.0030 0.0065
3.0 0.0025 0.0025 0.0029 0.0062
Case 3a
Strong correlation
No ZI
1.1 0.0027 0.0017 0.0013 0.0032
1.2 0.0129 0.0132 0.0052 0.0135
1.3 0.0105 0.0106 0.0059 0.0144
1.4 0.0077 0.0076 0.0054 0.0122
1.5 0.0059 0.0061 0.0049 0.0110
1.7 0.0040 0.0041 0.0045 0.0101
1.9 0.0040 0.0038 0.0032 0.0081
2.0 0.0030 0.0029 0.0029 0.0073
2.5 0.0022 0.0021 0.0030 0.0064
3.0 0.0015 0.0017 0.0026 0.0058
Case 3b
Strong correlation
30% ZI
1.1 0.0000 0.0000 0.0000 0.0074
1.2 0.0102 0.0106 0.0061 0.0153
1.3 0.0100 0.0100 0.0051 0.0120
1.4 0.0076 0.0075 0.0044 0.0115
1.5 0.0059 0.0058 0.0044 0.0109
1.7 0.0043 0.0041 0.0039 0.0094
1.9 0.0031 0.0032 0.0036 0.0082
2.0 0.0031 0.0031 0.0033 0.0078
2.5 0.0024 0.0023 0.0026 0.0063
3.0 0.0015 0.0015 0.0026 0.0059
@@
Case $\lambda$
DP Poisson
local-global
DP ZIP
local-global
DP Hu et al.
DP Hu et al.
posterior quantile optim
Case 1a
Weak correlation
No ZI
1.1 0.0005 0.0005 0.0001 0.0003
1.2 0.0296 0.0292 0.0166 0.0201
1.3 0.1080 0.1080 0.0548 0.0629
1.4 0.1678 0.1678 0.0979 0.1087
1.5 0.2143 0.2147 0.1409 0.1533
1.7 0.2854 0.2851 0.2130 0.2274
1.9 0.3404 0.3407 0.2738 0.2895
2.0 0.3647 0.3648 0.3002 0.3164
2.5 0.4584 0.4588 0.4061 0.4231
3.0 0.5245 0.5252 0.4818 0.4988
Case 1b
Weak correlation
30% ZI
1.1 0.0006 0.0004 0.0001 0.0003
1.2 0.0296 0.0291 0.0167 0.0203
1.3 0.1090 0.1086 0.0545 0.0625
1.4 0.1694 0.1700 0.0995 0.1102
1.5 0.2170 0.2165 0.1423 0.1546
1.7 0.2847 0.2849 0.2129 0.2269
1.9 0.3399 0.3396 0.2728 0.2883
2.0 0.3639 0.3639 0.2999 0.3152
2.5 0.4571 0.4575 0.4048 0.4211
3.0 0.5256 0.5261 0.4813 0.4998
Case 2a
Moderate correlation
No ZI
1.1 0.0027 0.0029 0.0003 0.0006
1.2 0.2210 0.2219 0.0455 0.0538
1.3 0.3322 0.3325 0.1104 0.1221
1.4 0.3989 0.3991 0.1677 0.1821
1.5 0.4410 0.4410 0.2193 0.2354
1.7 0.5011 0.5014 0.3000 0.3158
1.9 0.5415 0.5417 0.3638 0.3810
2.0 0.5572 0.5571 0.3905 0.4078
2.5 0.6167 0.6172 0.4976 0.5147
3.0 0.6597 0.6599 0.5667 0.5826
Case 2b
Moderate correlation
30% ZI
1.1 0.0028 0.0025 0.0004 0.0007
1.2 0.2187 0.2202 0.0454 0.0543
1.3 0.3323 0.3329 0.1098 0.1215
1.4 0.3977 0.3977 0.1666 0.1809
1.5 0.4411 0.4412 0.2202 0.2350
1.7 0.5002 0.5002 0.2999 0.3158
1.9 0.5408 0.5408 0.3632 0.3801
2.0 0.5571 0.5572 0.3914 0.4085
2.5 0.6177 0.6180 0.4974 0.5141
3.0 0.6611 0.6613 0.5674 0.5835
Case 3a
Strong correlation
No ZI
1.1 0.0053 0.0055 0.0004 0.0009
1.2 0.3080 0.3092 0.0480 0.0583
1.3 0.4558 0.4565 0.1219 0.1360
1.4 0.5480 0.5486 0.1875 0.2042
1.5 0.6090 0.6094 0.2466 0.2647
1.7 0.6805 0.6806 0.3412 0.3601
1.9 0.7177 0.7176 0.4124 0.4316
2.0 0.7290 0.7290 0.4449 0.4636
2.5 0.7675 0.7676 0.5590 0.5768
3.0 0.7959 0.7961 0.6307 0.6470
Case 3b
Strong correlation
30% ZI
1.1 0.0039 0.0044 0.0004 0.0007
1.2 0.3061 0.3076 0.0492 0.0593
1.3 0.4517 0.4530 0.1213 0.1356
1.4 0.5458 0.5462 0.1876 0.2041
1.5 0.6058 0.6061 0.2461 0.2644
1.7 0.6804 0.6805 0.3397 0.3588
1.9 0.7149 0.7150 0.4131 0.4330
2.0 0.7275 0.7276 0.4441 0.4636
2.5 0.7669 0.7670 0.5584 0.5761
3.0 0.7960 0.7962 0.6299 0.6462
@@
Case $\lambda$
DP Poisson
local-global
DP ZIP
local-global
DP Hu et al.
DP Hu et al.
posterior quantile optim
Case 1a
Weak correlation
No ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0000 0.0000 0.0000 0.0000
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 1b
Weak correlation
30% ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0000 0.0000 0.0000 0.0000
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 2a
Moderate correlation
No ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0000 0.0000 0.0000 0.0000
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 2b
Moderate correlation
30% ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0000 0.0000 0.0000 0.0000
1.4 0.0000 0.0000 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 3a
Strong correlation
No ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0001 0.0001 0.0000 0.0000
1.3 0.0001 0.0001 0.0000 0.0000
1.4 0.0001 0.0001 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
Case 3b
Strong correlation
30% ZI
1.1 0.0000 0.0000 0.0000 0.0000
1.2 0.0000 0.0000 0.0000 0.0000
1.3 0.0001 0.0001 0.0000 0.0000
1.4 0.0001 0.0001 0.0000 0.0000
1.5 0.0000 0.0000 0.0000 0.0000
1.7 0.0000 0.0000 0.0000 0.0000
1.9 0.0000 0.0000 0.0000 0.0000
2.0 0.0000 0.0000 0.0000 0.0000
2.5 0.0000 0.0000 0.0000 0.0000
3.0 0.0000 0.0000 0.0000 0.0000
@@
Case $\lambda$
DP Poisson
local-global
DP ZIP
local-global
DP Hu et al.
DP Hu et al.
posterior quantile optim
Case 1a
Weak correlation
No ZI
1.1 0.0009 0.0009 0.0003 0.0006
1.2 0.0564 0.0556 0.0323 0.0390
1.3 0.1927 0.1927 0.1032 0.1173
1.4 0.2857 0.2858 0.1773 0.1948
1.5 0.3513 0.3519 0.2459 0.2644
1.7 0.4423 0.4420 0.3498 0.3688
1.9 0.5062 0.5065 0.4286 0.4472
2.0 0.5328 0.5329 0.4605 0.4788
2.5 0.6269 0.6272 0.5762 0.5925
3.0 0.6864 0.6869 0.6489 0.6635
Case 1b
Weak correlation
30% ZI
1.1 0.0011 0.0009 0.0003 0.0007
1.2 0.0564 0.0554 0.0326 0.0394
1.3 0.1946 0.1939 0.1025 0.1168
1.4 0.2881 0.2890 0.1799 0.1974
1.5 0.3551 0.3544 0.2481 0.2665
1.7 0.4416 0.4418 0.3498 0.3681
1.9 0.5056 0.5053 0.4273 0.4456
2.0 0.5319 0.5319 0.4600 0.4773
2.5 0.6257 0.6261 0.5749 0.5907
3.0 0.6875 0.6879 0.6486 0.6646
Case 2a
Moderate correlation
No ZI
1.1 0.0049 0.0052 0.0006 0.0011
1.2 0.3585 0.3596 0.0864 0.1014
1.3 0.4960 0.4964 0.1981 0.2166
1.4 0.5683 0.5684 0.2863 0.3067
1.5 0.6099 0.6099 0.3586 0.3795
1.7 0.6661 0.6662 0.4605 0.4784
1.9 0.7010 0.7012 0.5324 0.5498
2.0 0.7144 0.7143 0.5606 0.5774
2.5 0.7616 0.7620 0.6633 0.6775
3.0 0.7938 0.7939 0.7223 0.7342
Case 2b
Moderate correlation
30% ZI
1.1 0.0052 0.0047 0.0008 0.0014
1.2 0.3552 0.3571 0.0863 0.1024
1.3 0.4959 0.4964 0.1972 0.2157
1.4 0.5672 0.5672 0.2848 0.3050
1.5 0.6103 0.6104 0.3600 0.3791
1.7 0.6652 0.6652 0.4604 0.4783
1.9 0.7006 0.7007 0.5317 0.5490
2.0 0.7144 0.7144 0.5614 0.5781
2.5 0.7626 0.7627 0.6631 0.6770
3.0 0.7948 0.7949 0.7227 0.7347
Case 3a
Strong correlation
No ZI
1.1 0.0094 0.0097 0.0009 0.0017
1.2 0.4665 0.4678 0.0911 0.1095
1.3 0.6224 0.6230 0.2167 0.2386
1.4 0.7050 0.7055 0.3150 0.3378
1.5 0.7545 0.7547 0.3947 0.4171
1.7 0.8081 0.8081 0.5078 0.5276
1.9 0.8339 0.8339 0.5830 0.6010
2.0 0.8420 0.8420 0.6149 0.6316
2.5 0.8674 0.8675 0.7160 0.7295
3.0 0.8856 0.8856 0.7725 0.7836
Case 3b
Strong correlation
30% ZI
1.1 0.0068 0.0077 0.0008 0.0014
1.2 0.4644 0.4661 0.0934 0.1113
1.3 0.6186 0.6198 0.2158 0.2380
1.4 0.7031 0.7034 0.3151 0.3378
1.5 0.7519 0.7521 0.3942 0.4167
1.7 0.8079 0.8080 0.5062 0.5262
1.9 0.8324 0.8324 0.5836 0.6023
2.0 0.8409 0.8410 0.6140 0.6315
2.5 0.8670 0.8671 0.7156 0.7290
3.0 0.8857 0.8858 0.7718 0.7829